The following equations describe an economy. (Think of C , I , G , etc., as being measured in
billions and i as a percentage; a 5 percent interest rate implies i = 5.)
C 5 0.8(1 2 t ) Y (P1)
t 5 0.25 (P2)
I 5 900 2 50 i (P3)
−−G 5 800 (P4)
L 5 0.25 Y 2 62.5 i (P5)
−−My−
P 5 500
4.* a. Show that a given change in the money stock has a larger effect on output the less
interest-sensitive is the demand for money. Use the formal analysis of Section 11-5.
b. How does the response of the interest rate to a change in the money stock depend on the
interest sensitivity of money demand?
1
Expert's answer
2021-08-16T10:00:36-0400

Question 4 a

a) Y=C+I+G=0.8(1-t)Y+900-50R+800=0.6Y+900-50R+800=0.8Y+1700.8-50R

Y-0.6Y=1700-50R

0.4Y=1700-50R

Y=4250-125R

Question 4b

The LM curve (liquidity-money) depicts the money market equilibrium when the demand for money (owing to the absolute cash liquidity property) equals the money supply. Because the interest rate influences the demand for money, the LM (Liquidity preference = Money supply) curve exists. Each point represents a mix of income and interest rates, maintaining monetary equilibrium.